# Operational Derivation of Born’s Rule from Causal Consistency in Generalized Probabilistic Theories

### v1

@@ 1 Introduction
Several derivations exist, but each comes with significant presuppositions. Gleason’s theorem [3] derives the Born rule from noncontextuality but assumes Hilbert-space structure from the outset and requires dimension at least three. Decision-theoretic approaches [4, 5] derive it through rationality arguments but rely on the controversial many-worlds interpretation and axioms that are not purely operational. Zurek’s envariance-based derivation [6] provides physical insights into environment-assisted invariance but still operates fully within the standard quantum formalism, taking the Hilbert-space machinery as given.
@@ 7.4 Limitations and Extensions
An important open direction is to identify the minimal subset of axioms needed to force $\Phi(p)=p$. For instance, it would be interesting to understand whether strong symmetry can be weakened, or whether some restricted form of purification suffices. Moreover, one could seek explicit quantitative bounds: how small can deviations from affinity be before they lead to experimentally detectable signaling given realistic noise and finite statistics?
@@ 8 Conclusion
We have established that Born’s rule emerges from causal consistency within the GPT framework. First, working operationally without presupposing Hilbert space structure, we showed that no superluminal signaling and normalization on sharp tests, together with operational affinity, force the rule $P(\phi|\psi)=\tau(\psi,\phi)$. Second, standard reconstruction theorems identify the GPT with complex quantum theory, for which $\tau(\psi,\phi)=|\langle\phi|\psi\rangle|^{2}$. Third, any non-affine deviation leads to superluminal signaling via steering, thereby explaining the universality and stability of quantum probabilities from information-theoretic principles.

### v2

@@ 1 Introduction
Several derivations exist, but each comes with significant presuppositions. Gleason’s theorem [3] derives the Born rule from noncontextuality but assumes Hilbert-space structure from the outset and requires dimension at least three. Decision-theoretic approaches [4, 5] derive it through rationality arguments but rely on the controversial many-worlds interpretation and axioms that are not purely operational. Zurek’s envariance-based derivation [6] provides physical insights into environment-assisted invariance but still operates fully within the standard quantum formalism, taking the Hilbert-space machinery as given. Relatedly, recent work has explored the status of the Born rule within abstract process-theoretic frameworks. In particular, Agrawal and Wilson [20] showed that any probabilistic process theory can be transformed, via suitable categorical quotient constructions, into an operationally equivalent theory satisfying a generalized Born rule. Their result clarifies the structural relationship between compositional rules and probability assignments in process theories.
@@ 7.4 Limitations and Extensions
An important open direction is to identify the minimal subset of axioms needed to force $\Phi(p)=p$. For instance, it would be interesting to understand whether strong symmetry can be weakened, or whether some restricted form of purification suffices. Moreover, one could seek explicit quantitative bounds: how small can deviations from affinity be before they lead to experimentally detectable signaling, given realistic noise and finite statistics?
@@ 8 Conclusion
We have established that Born’s rule emerges as a consequence of causal consistency within the GPT framework. Working operationally and without presupposing Hilbert-space structure, we showed that no superluminal signaling, together with normalization on sharp tests and operational affinity under mixing, uniquely fixes the probability assignment to $P(\phi|\psi)=\tau(\psi,\phi)$. Standard reconstruction theorems then recover complex quantum theory, where $\tau(\psi,\phi)=|\langle\phi|\psi\rangle|^{2}$.
Crucially, any non-affine deviation from this rule generically enables superluminal signaling via steering, identifying the Born rule as a causal fixed point among admissible probabilistic laws. In this sense, the universality and robustness of quantum probabilities are not mere artifacts of Hilbert-space geometry, but reflect deep information-theoretic and causal constraints.